# Picosecond laser test unit for photosensor characterization

### v1

@@ 1 Introduction
Modern low-background neutrino and dark matter experiments rely on large arrays of photosensors to detect faint scintillation and Cherenkov light. Accurate knowledge of the single photoelectron (SPE) response, gain, and timing is required to control systematic uncertainties in reconstructed energy, position, and event topology. Photomultiplier tubes (PMTs) are widely used in liquid scintillator and Cherenkov detectors because they combine SPE sensitivity with nanosecond-scale timing. Dedicated test stands have supported PMT qualification and calibration for large experiments such as for Borexino [BRIGATTI2005521, PMT_SmiLomRan2004], Double Chooz [QualiPMTs_DC_2011, TTandCh_SPE_R7081_PMTs_2012], XENON1T [QualiPMTs_XENON_2017] and JUNO [JUNO], typically emphasizing high-throughput testing, robust calibration procedures, and long-term stability.
In parallel, more compact and flexible laboratory setups are needed for detailed PMT and photosensor studies in R&D environments. Comprehensive photosensor studies for large liquid scintillator detectors, for example in the context of the proposed LENA observatory, have highlighted this need [PhDTippmann]. Building on this experience, we employ a compact, table-top bench with picosecond laser excitation, controlled illumination, and synchronized data acquisition (DAQ) for photosensor characterization, and demonstrate its performance with PMTs.
@@
While many aspects of PMT characterization are well established, this work provides a concise, practice-oriented reference that combines a compact setup with waveform-based analysis procedures applicable to a wide range of PMT types. In addition to the standard charge and timing observables, a model-independent method is introduced to quantify the double photoelectron (DPE) contribution directly from the pulse charge spectrum. The characterization procedures developed here provide the basis for describing the PMT responses in our liquid scintillator (LS) R&D program and are used to verify and validate the performance of the LS test setup that underlies the measurements reported in [WbLS, SlowLS, TAUP2023].
@@ 2 Experimental setup
The effective temporal width of the laser pulses was verified using a KETEK PM3325-WB SiPM placed 3.3 cm from the fiber output, operated at 29.7 V. The time difference between SiPM signals and the laser trigger follows a Gaussian distribution with FWHM $(12.4\pm 0.1)$ ps [PhD_MRStock2025], well below the PMT TTS values studied here ($\gtrsim 180$ ps), so no deconvolution of transit time spectra is required.
@@
The PMT is mounted either in a light-tight, blackened aluminum box (approximately $70\times 54\times 40\,\mathrm{cm}^{3}$ in size) or in a Binder MKT 115 climate chamber. Both provide optical isolation and act as Faraday cages. The fiber and cables enter through light-tight feedthroughs, and the fiber is positioned about 35 cm from the photocathode, as shown in figure 2, ensuring full coverage of the active area. Under far-field geometry, the laser spot size increases approximately linearly with the fiber-to-photocathode distance. Although the irradiance is not completely uniform across the photocathode, this geometry is sufficient for SPE-level characterization of 3-inch PMTs. A detailed study of the illumination profile can be found in [PhD_MRStock2025]. For temperature studies, the PMT, fiber and signal cables are placed inside the climate chamber, operated within $50^{\circ}$C to $20^{\circ}$C. After each temperature change, data taking begins only after sufficient equilibration time of all components.
@@
The PMTs are powered by a CAEN N1470 high-voltage supply with 1 V adjustment precision. A Rohde & Schwarz RTO2004 oscilloscope generates two synchronized logic signals. A TTL pulse triggers the laser, while a fast-negative NIM pulse is used to trigger data acquisition. The NIM trigger can be phase-shifted so that both the PMT pulse and trigger lie within the selected acquisition time window. The laser repetition rate is reduced sufficiently such that one trigger occurs within each ADC window. The timing precision of the waveform generator was verified, producing a narrow dominant Gaussian distribution with FWHM $(8.4\pm 0.2)$ ps and an overall distribution width of 33 ps [PhD_MRStock2025], both negligible compared to PMT transit time spreads.
@@
Data are stored in binary format and converted into ROOT [ROOT] files using a dedicated C++ software framework originally developed for the positronium lifetime determination experiment [SCHWARZ201964]. The software reconstructs waveforms and extracts parameters such as baseline parameters, pulse height, integrated charge, and timing information. A graphical interface supports validation and inspection of individual events.
@@ 3 Pulse characterization
    Trace integral: area under the entire baseline-corrected waveform, providing a measure of the total collected charge. This quantity includes noise outside the pulse region and thus depends on the acquisition window length.
@@
    Pulse integral: area under the pulse above 10 % of its pulse height relative to the baseline. This definition excludes baseline contributions outside the pulse and is therefore less sensitive to the chosen ADC window.
@@ 4 Data analysis and methods
The following analysis procedures are illustrated using example data from a 3-inch 9821B PMT operated at 1800 V supply voltage. All results shown here are representative of the general methodology applied to all tubes. Each measurement typically contains more than one million events, allowing precise determination of the distributions of relevant pulse descriptors. Histogram bin widths are chosen as small as permitted by the ADC sampling granularity, maximizing resolution while avoiding digitization artifacts. All analysis is performed in Python 3.8 [python] using numpy 1.24.3 [numpy]scipy 1.10.1 [scipy], matplotlib 3.72 [matplotlib], uproot 5.3.13 [uproot], iminuit 2.29.1 [iminuit], SymPy 1.13.3 [Sympy] and pandas 2.0.3 [pandas]. The basic event selection and preprocessing steps are as follows:
@@
    The baseline mean and baseline standard deviation spectra are modeled with one to three Gaussian components using an extended binned maximum likelihood fit (via iminuit [iminuit]). The preferred model is chosen via the reduced $\chi^{2}$, and $^{\prime}3\,\sigma^{\prime}$ selection cuts are typically applied to suppress outliers.
@@
    The pulse number is constrained to at most one identified pulse for noise studies of the trace charge spectrum and to exactly one otherwise.
@@
    The pulse height distribution is approximated using a kernel density estimation (KDE) implemented in scipy.stats.gaussian_kde [scipy]. This method provides a smooth, binning-independent representation of the distribution. Its bandwidth is tuned to avoid artificial structure while retaining physical features. Depending on the pulse identifier requirement, either a distinct valley above the noise peak is visible or the noise population is fully suppressed. If the valley is visible, its position is taken from the KDE and used as the lower pulse height threshold to reject noise events. If the valley is absent, a fixed lower threshold is used, e.g., the voltage at which the KDE reaches 50 % of its maximum. Saturated events close to the ADC voltage limit are identified and removed. Except for dedicated studies of the trace charge spectrum, both the pulse height noise and saturated components are excluded from the analysis.
After these basic data quality selections, the data are ready for extraction of trace charge spectra, pulse charge spectra, and transit time spectra [PhD_MRStock2025]. The following subsections describe these procedures and corresponding statistical models.
@@ 4.1 Trace charge spectrum
Trace integrals are used to select SPE events, extract the SPE mean charge $\mu_{\text{SPE}}$, and determine the PMT gain $G$. The integrals are recorded in mVns and converted to charge (pC) using $q={U\times\Delta t}/{R}$ with the back-termination resistance $R=50\,\Omega$. Bins containing saturated events are excluded from the fit by scanning downward from the maximum charge bin. To describe the trace charge distribution, we adopt and extend the model of [Diwan2020]. At low-light levels, the number of photoelectrons (PEs) per trigger $k$ follows a Poisson distribution with mean $\lambda$. The probability of observing $k$ PEs is then, $P(k\mid\lambda)={\lambda^{k}\,\mathrm{e}^{-\lambda}}/{k!}$, which serves as the statistical weighting factor for each $k$ PE component.
@@
where $q$ is the charge variable, $q_{0}$ is the mean pedestal charge, and $\sigma_{0}>0$ is the noise standard deviation [Diwan2020]. The second background component comes from discrete processes (thermionic dark pulses or under-amplified signals [QualiPMTs_DC_2011, QualiPMTs_XENON_2017, Diwan2020]). These processes include PEs missing the first dynode and PE emission from focusing electrodes or dynodes, empirically described by
@@
In the SPE case ($k=1$), PEs collected at the first dynode induce the emission of secondary electrons, which multiply further through the dynode structure. The amplification factor of this process is known as gain $G$, whose fluctuation is mainly dominated by the first dynode. Under the assumption of a sufficiently large first-dynode secondary-emission coefficient, e.g., >4>4, and a collection coefficient near unity, the SPE charge distribution can be modeled as a Gaussian distribution $G_{\text{PDF}}(q\,|\,q_{1},\sigma_{1})$ with mean charge $q_{1}$ and standard deviation $\sigma_{1}$ [Diwan2020].
@@
However, following [RDossi2000], the signal components are here truncated below the pedestal mean $q_{0}$ using the Heaviside step function $\Theta(q-q_{0})$ to better describe the data. Therefore, a Heaviside modified Gaussian distribution is defined as
@@
Table 1: Parameters of the trace charge spectrum model (see Eq. 4.1).
Fits are performed with an extended binned maximum likelihood method using the iminuit library [iminuit]. Figure 4 shows a fit of the model, lists individual components and best-fit parameters. The lower panel highlights the residuals and indicates the data range used for fitting (colored in blue), excluding saturated events. The probability of dark pulses, $w_{0}\approx 2\,\%$, is relatively small. The mean number of photoelectrons, here $\lambda=0.06$, can be used to derive the occupancy $O_{\text{PMT}}=1-\exp(-\lambda)=6\,\%$, implying that out of every hundred laser pulses, only six are detected by the PMT and confirming the low-intensity light [PMT_DEAP3600_2019]. The peak in the SPE regime $P$ is identified as the local maximum of the fitted charge spectrum above the pedestal. The valley $V$ is defined as the local minimum between pedestal and $P$. Uncertainties of both peak and valley are determined from the $1\sigma$ uncertainty band of the fit, which was derived from the covariance matrix of the fitted parameters. By this, the peak-to-valley ratio (P/V) can be determined, quantifying the separation power between noise and signal [QualiPMTs_XENON_2017].
@@
This corrected mean value is highlighted in green in figure 4. Similarly, the true standard deviation $\sigma_{\text{SPE}}$ is computed. Uncertainties of both values are determined using uncertainty propagation, based on the best-fit parameters and their statistical uncertainties. The partial derivatives required for this propagation were numerically computed using the Sympy library [Sympy].
The gain is calculated by $G=\mu_{\text{SPE}}/e$ [PMT_Bellamy_1994], where $e$ is the electric charge. Here, the gain is $G=5\times 10^{6}$ at 1800 V, which is in agreement with the datasheet [9821B_datasheet].
@@ 4.2 Pulse charge spectrum
Figure 5: Example pulse charge spectrum at SPE intensity for a 9821B PMT operated at 1800 V. The gold curve is a Gaussian KDE and the green marker indicates the SPE mean charge. Vertical red dashed lines denote the SPE selection window. The cyan dash-dotted curve is a DPE model fitted to the high-charge tail to estimate DPE contamination within the SPE window.
Assuming most entries originate from SPE pulses, the histogram is self-convolved to determine the DPE distribution’s position and shape. This self-convolution approach provides a model-independent estimate of the DPE component, avoiding assumptions about the underlying SPE shape and requiring only the measured pulse charge distribution. Since the DPE shape is fixed, the number of DPE events $N_{\text{DPE}}$ is the sole free parameter in the following fit. The DPE fit is first performed on bins with charge greater than the DPE peak. The resulting DPE component is compared with the data, and bins in which it contributes $75\,\%$ of the content are selected. A second fit restricted to these bins mitigates overestimation and yields more reliable results. In figure 5, the solid turquoise line indicates the data range used for the fit, while the full DPE component is depicted with a dash-dotted line. The SPE mean $\mu_{\text{SPE}}$, standard deviation $\sigma_{\text{SPE}}$, and the uncertainty of the SPE mean charge are estimated directly from the data excluded from the DPE fit. The SPE mean from pulse integrals, here $0.85\,\text{pC}$, is systematically lower than that from trace integrals, here $0.99\,\text{pC}$, because the pulse integral method does not fully capture the pulse area (section 3). The pulse charge spectrum is approximated via KDE. Lower and upper charge cuts for SPE selection can be set using (i) the charge where the KDE falls to $50\,\%$ of its maximum and (ii) the DPE peak, respectively. Here, the resulting DPE contamination inside the chosen SPE interval is $\sim 2\,\%$, comparable to the $\sim 1.4\,\%$ obtained from the trace charge model fit. Alternatively, data selection intervals can be defined in PE units, e.g., 0.50.51.51.5 PE.

### v2

@@ 1 Introduction
Modern low-background neutrino and dark matter experiments rely on large arrays of photosensors to detect faint scintillation and Cherenkov light. Accurate knowledge of the single photoelectron (SPE) response, gain, and timing is required to control systematic uncertainties in reconstructed energy, position, and event topology. Photomultiplier tubes (PMTs) are widely used in liquid scintillator and Cherenkov detectors because they combine SPE sensitivity with nanosecond-scale timing. Dedicated test stands have supported PMT qualification and calibration for large experiments such as for Borexino [2, 25], Double Chooz [5, 10], XENON1T [26] and JUNO [1], typically emphasizing high-throughput testing, robust calibration procedures, and long-term stability.
In parallel, more compact and flexible laboratory setups are needed for detailed PMT and photosensor studies in R&D environments. Comprehensive photosensor studies for large liquid scintillator detectors, for example in the context of the proposed LENA observatory, have highlighted this need [20]. Building on this experience, we employ a compact, table-top bench with picosecond laser excitation, controlled illumination, and synchronized data acquisition (DAQ) for photosensor characterization, and demonstrate its performance with PMTs.
@@
While many aspects of PMT characterization are well established, this work provides a concise, practice-oriented reference that combines a compact setup with waveform-based analysis procedures applicable to a wide range of PMT types. In addition to the standard charge and timing observables, a model-independent method is introduced to quantify the double photoelectron (DPE) contribution directly from the pulse charge spectrum. The characterization procedures developed here provide the basis for describing the PMT responses in our liquid scintillator (LS) R&D program and are used to verify and validate the performance of the LS test setup that underlies the measurements reported in [15, 14, 21].
@@ 2 Experimental setup
The effective temporal width of the laser pulses was verified using a KETEK PM3325-WB SiPM placed 3.3 cm from the fiber output, operated at 29.7 V. The time difference between SiPM signals and the laser trigger follows a Gaussian distribution with FWHM $(12.4\pm 0.1)$ ps [22], well below the PMT TTS values studied here ($\gtrsim 180$ ps), so no deconvolution of transit time spectra is required.
@@
The PMT is mounted either in a light-tight, blackened aluminum box (approximately $70\times 54\times 40\,\mathrm{cm}^{3}$ in size) or in a Binder MKT 115 climate chamber. Both provide optical isolation and act as Faraday cages. The fiber and cables enter through light-tight feedthroughs, and the fiber is positioned about 35 cm from the photocathode, as shown in figure 2, ensuring full coverage of the active area. Under far-field geometry, the laser spot size increases approximately linearly with the fiber-to-photocathode distance. Although the irradiance is not completely uniform across the photocathode, this geometry is sufficient for SPE-level characterization of 3-inch PMTs. A detailed study of the illumination profile can be found in [22]. For temperature studies, the PMT, fiber and signal cables are placed inside the climate chamber, operated within $50^{\circ}$C to $20^{\circ}$C. After each temperature change, data taking begins only after sufficient equilibration time of all components.
@@
The PMTs are powered by a CAEN N1470 high-voltage supply with 1 V adjustment precision. A Rohde & Schwarz RTO2004 oscilloscope generates two synchronized logic signals. A TTL pulse triggers the laser, while a fast-negative NIM pulse triggers data acquisition. The NIM trigger can be phase-shifted so that both the PMT pulse and trigger lie within the selected DAQ time window.
The laser repetition rate is typically 20 kHz or lower (e.g., 10 kHz), corresponding to pulse spacings of 50–100 µs. The DAQ time window is 100 ns, which is more than five orders of magnitude shorter than the interval between consecutive laser pulses. Consequently, each recorded waveform contains the PMT response to at most one laser pulse, preventing overlap of independent events within a single acquisition time window.
The timing precision of the waveform generator was verified by generating synchronized  TTL and NIM logic signals from the RTO2004 oscilloscope and recording the waveforms at 10 kHz using the internal 10 GS/s ADC, with interpolation applied to improve resolution. The resulting time difference distribution exhibits a dominant Gaussian component with FWHM $(8.6\pm 0.2)$ ps and an overall width of 33 ps [22], both negligible compared to typical PMT transit time spreads.
@@
Data are stored in binary format and converted into ROOT [11] files using a dedicated C++ software framework originally developed for the positronium lifetime determination experiment [23]. The software reconstructs waveforms and extracts parameters such as baseline parameters, pulse height, integrated charge, and timing information. A graphical interface supports validation and inspection of individual events.
@@ 3 Pulse characterization
    Trace integral: area under the entire baseline-corrected waveform, providing a measure of the total collected charge. This quantity includes noise outside the pulse region and thus depends on the DAQ time window.
@@
    Pulse integral: area under the pulse above 10 % of its pulse height relative to the baseline. This definition excludes baseline contributions outside the pulse and is therefore less sensitive to the chosen DAQ time window.
@@ 4 Data analysis and methods
The following analysis procedures are illustrated using example data from a 3-inch 9821B PMT operated at 1800 V supply voltage. All results shown here are representative of the general methodology applied to all tubes. Each measurement typically contains more than one million events, allowing precise determination of the distributions of relevant pulse descriptors. Histogram bin widths are chosen as small as permitted by the ADC sampling granularity, maximizing resolution while avoiding digitization artifacts. All analysis is performed in Python 3.8 [12] using numpy 1.24.3 [6]scipy 1.10.1 [27], matplotlib 3.72 [18], uproot 5.3.13 [19], iminuit 2.29.1 [13], SymPy 1.13.3 [3] and pandas 2.0.3 [30]. The basic event selection and preprocessing steps are as follows:
@@
    The baseline mean and baseline standard deviation spectra are modeled with one to three Gaussian components using an extended binned maximum likelihood fit (via iminuit [13]). The preferred model is chosen via the reduced $\chi^{2}$, and $^{\prime}3\,\sigma^{\prime}$ selection cuts are typically applied to suppress outliers.
@@
    For noise studies of the trace charge spectrum, waveforms with zero or one identified pulse are retained. In all other cases, such as transit time spectrum studies, exactly one pulse per waveform is retained.
@@
    The pulse height distribution is approximated using a Gaussian kernel density estimation (KDE) implemented in scipy.stats.gaussian_kde [27]. This KDE method provides a smooth, binning-independent representation of the distribution. Its bandwidth is tuned to avoid artificial structure while retaining physical features. Depending on the pulse identifier requirement, either a distinct valley above the noise peak is visible or the noise population is fully suppressed. If the valley is visible, its position is taken from the KDE and used as the lower pulse height threshold to reject noise events. If the valley is absent, a fixed lower threshold is used, e.g., the voltage at which the KDE reaches 50 % of its maximum. Saturated events close to the ADC voltage limit are identified and removed. Except for dedicated studies of the trace charge spectrum, both the pulse height noise and saturated components are excluded from the analysis.
After these basic data quality selections, the data are ready for extraction of trace charge spectra, pulse charge spectra, and transit time spectra [22]. The following subsections describe these procedures and corresponding statistical models.
@@ 4.1 Trace charge spectrum
Trace integrals are used to select SPE events, extract the SPE mean charge $\mu_{\text{SPE}}$, and determine the PMT gain $G$. The integrals are recorded in mVns and converted to charge (pC) using $q={U\times\Delta t}/{R}$ with the back-termination resistance $R=50\,\Omega$. Bins containing saturated events are excluded from the fit by scanning downward from the maximum charge bin. To describe the trace charge distribution, we adopt and extend the model of [24]. At low-light levels, the number of photoelectrons (PEs) per trigger $k$ follows a Poisson distribution with mean $\lambda$. The probability of observing $k$ PEs is then, $P(k\mid\lambda)={\lambda^{k}\,\mathrm{e}^{-\lambda}}/{k!}$, which serves as the statistical weighting factor for each $k$ PE component.
@@
where $q$ is the charge variable, $q_{0}$ is the mean pedestal charge, and $\sigma_{0}>0$ is the noise standard deviation [24]. The second background component comes from discrete processes (thermionic dark pulses or under-amplified signals [5, 26, 24]). These processes include PEs missing the first dynode and PE emission from focusing electrodes or dynodes, empirically described by
@@
In the SPE case ($k=1$), PEs collected at the first dynode induce the emission of secondary electrons, which multiply further through the dynode structure. The amplification factor of this process is known as gain $G$, whose fluctuation is mainly dominated by the first dynode. Under the assumption of a sufficiently large first-dynode secondary-emission coefficient, e.g., >4>4, and a collection coefficient near unity, the SPE charge distribution can be modeled as a Gaussian distribution $G_{\text{PDF}}(q\,|\,q_{1},\sigma_{1})$ with mean charge $q_{1}$ and standard deviation $\sigma_{1}$ [24].
@@
However, following [29], the signal components are here truncated below the pedestal mean $q_{0}$ using the Heaviside step function $\Theta(q-q_{0})$ to better describe the data. Therefore, a Heaviside modified Gaussian distribution is defined as
@@
Table 1: Parameters of the trace charge spectrum model (see equation 4.1).
Fits are performed with an extended binned maximum likelihood method using the iminuit library [13]. Figure 4 shows a fit of the model, lists individual components and best-fit parameters. The lower panel highlights the residuals and indicates the data range used for fitting (colored in blue), excluding saturated events. The probability of dark pulses, $w_{0}\approx 2\,\%$, is relatively small. The mean number of photoelectrons, here $\lambda=0.06$, can be used to derive the occupancy $O_{\text{PMT}}=1-\exp(-\lambda)=6\,\%$, implying that out of every hundred laser pulses, only six are detected by the PMT and confirming the low-intensity light [28]. The peak in the SPE regime $P$ is identified as the local maximum of the fitted charge spectrum above the pedestal. The valley $V$ is defined as the local minimum between pedestal and $P$. Uncertainties of both peak and valley are determined from the $1\sigma$ uncertainty band of the fit, which was derived from the covariance matrix of the fitted parameters. By this, the peak-to-valley ratio (P/V) can be determined, quantifying the separation power between noise and signal [26].
@@
This corrected mean value is highlighted in green in figure 4. Similarly, the true standard deviation $\sigma_{\text{SPE}}$ is computed. Uncertainties of both values are determined using uncertainty propagation, based on the best-fit parameters and their statistical uncertainties. The partial derivatives required for this propagation were numerically computed using the Sympy library [3].
The gain is calculated by $G=\mu_{\text{SPE}}/e$ [7], where $e$ is the electric charge. Here, the gain is $G=5\times 10^{6}$ at 1800 V, which is in agreement with the datasheet [8].
@@ 4.2 Pulse charge spectrum
Figure 5: Example pulse charge spectrum at SPE intensity for a 9821B PMT operated at 1800 V. The gold curve is a Gaussian KDE, and the green marker indicates the SPE mean charge. Vertical red dashed lines denote the SPE selection window. The cyan dash-dotted curve is a DPE model fitted to the high-charge tail to estimate DPE contamination within the SPE window, while the solid cyan line indicates the data range used for the fit.
Assuming that most entries originate from SPE pulses, the histogram is self-convolved to determine the position and shape of the DPE distribution. This self-convolution approach provides a model-independent estimate of the DPE component, avoiding assumptions about the underlying SPE shape and requiring only the measured pulse charge distribution. Since the DPE shape is fixed, the number of DPE events $N_{\text{DPE}}$ is the sole free parameter in the fit. The DPE fit is first performed on bins with charge greater than the DPE peak. The resulting DPE component is compared with the data, and bins in which it contributes $75\,\%$ of the content are selected. A second fit restricted to these bins mitigates overestimation and improves the stability of the DPE estimation. In figure 5, the solid cyan line indicates the data range used for the fit, while the full DPE component is depicted with a dash-dotted line.
The SPE mean $\mu_{\text{SPE}}$, standard deviation $\sigma_{\text{SPE}}$, and the uncertainty of the SPE mean charge are estimated directly from the data excluded from the DPE fit. By excluding data clearly associated with DPE events, this approach provides a slightly improved estimate compared to simply taking the mean of the entire distribution. The SPE mean from pulse integrals, here $0.85\,\text{pC}$, is systematically lower than that from trace integrals, here $0.99\,\text{pC}$, because the pulse integral method does not fully capture the pulse area (Section 3). The pulse charge spectrum is approximated via KDE. Lower and upper charge cuts for SPE selection can be set using (i) the charge where the KDE falls to $50\,\%$ of its maximum and (ii) the DPE peak, respectively. Here, the resulting DPE contamination inside the chosen SPE interval is $\sim 2\,\%$, comparable to the $\sim 1.4\,\%$ obtained from the trace charge model fit. Alternatively, data selection intervals can be defined in PE units, e.g., 0.50.51.51.5 PE.